OpenStudy (theoreo):

Find where the points on the given curve on concave up in interval notation. I've managed to get the derivatives of the curve, and I think I've got the critical points, but I'm having trouble figuring out if the graph is concave up at all, and if it's not, I forgot how to state that in interval notation. I'll put the equations in the next post.

OpenStudy (anonymous):

Have you taken the second derivative?

OpenStudy (theoreo):

Yeah I've got the first and second. I just wanted to use the equation thing to post them because they're kinda messy.

OpenStudy (theoreo):

\[\frac{ dy }{ dx } = \frac{ (\sin(t) }{ 2\sin(2t) }\] \[\frac{ d^2dy }{ dx^2 } = \frac{ 2\sin(2t)*\cos(t)-\sin(t)*4\cos(2t) }{ 2\sin(2t)^2*-\sin(2t) }\]

OpenStudy (theoreo):

Forgot to add, the boundary is from 0 < t < pi. So going off the first derivative the critical points should be 0 and pi right?

OpenStudy (anonymous):

For the the second derivative, you used f'(x)g(x)-g'(x)f(x) / (g(x))^2?

OpenStudy (theoreo):

Yeah and I verified it's correct.

OpenStudy (anonymous):

And yes, those critical points seem right.

OpenStudy (anonymous):

You just have to find the points of inflection and plug them in then.

OpenStudy (anonymous):

You want to find concavity, right?

OpenStudy (theoreo):

Yes. I'm looking for where the curve is concave up.

OpenStudy (anonymous):

Are you having trouble finding the inflection points then?

OpenStudy (theoreo):

I don't think so. I think I've managed to solve the second derivative down so that the inflection points are also 0 and pi. My problem is, I'm supposed to run a value between the critical points to determine concavity right? Well most of the values I tried (pi/2, pi/4 etc) just caused the denominator to zero out. I managed to find a couple of values that actually gave me a result, but they were all negative. So I think the curve is concave down over the entire interval. If that's correct, I don't know how to state that in interval notation.

OpenStudy (anonymous):

Are you sure your second derivative is right? I'm not entirely certain why you multiplied (2sin(2t))^2 by -sin(2t)

OpenStudy (theoreo):

Yes, I'm using an online homework submission program, and it's saying that second derivative is correct.

OpenStudy (anonymous):

I'll see what I get for inflection points then first.

OpenStudy (anonymous):

I'm getting something different, but it doesn't seem to really work either.

OpenStudy (anonymous):

Oh, never mind, I messed up a negative.

OpenStudy (anonymous):

You said you got 0 and pi?

OpenStudy (anonymous):

Do you know if they're right?

OpenStudy (theoreo):

Yes.

OpenStudy (theoreo):

No, I'm not sure if those are right. On second thought, it seems odd that the critical points would also be inflection points. Is that possible?

OpenStudy (anonymous):

No, that's possible. I think you're right though, I got those numbers too once I plugged in the identity right.

OpenStudy (anonymous):

You were right from the start, unless you plugged something in wrong.

OpenStudy (theoreo):

Ok. So if the curve isn't concave up along the interval, is there a way to say that in interval notation?

OpenStudy (anonymous):

Isn't it positive from 0 to pi though?

OpenStudy (theoreo):

I put pi/4 into my graphing calculator and the value came back as -1 in the equation for the second derivative. Unless I entered it wrong, I'll check.

OpenStudy (anonymous):

I plugged in 1.5 and it came out positive.

OpenStudy (anonymous):

and then pi/4 comes out negative.

OpenStudy (theoreo):

I double checked the equation I put in and I'm getting a negative value for 1.5 and pi/4. Even if that's wrong though, if 1.5 is positive and pi/4 is negative... That doesn't track with the critical points.

OpenStudy (anonymous):

Yeah. I'm probably doing something stupid. @sourwing can probably help you better.

OpenStudy (theoreo):

I doubt it's you. This question seems pretty ridiculous. I've run it by a tutor it seems to have stumped them as well.

OpenStudy (anonymous):

what are the y(t) and x(t). There is something not quite right in your second derivative. But I need y(t) and x(t)

OpenStudy (theoreo):

x = cos(2t) y = cos(t)

OpenStudy (theoreo):

The second derivative should be right unless there's some kind of error with the program though.

OpenStudy (anonymous):

ok, did you do d/dt (-sint / (-2sin(2t)) ------------------------- d/dt (cos(2t)) ??

OpenStudy (theoreo):

Yes and I ended up with the second derivative I posted.

OpenStudy (anonymous):

ok, did you get (1/4) (sect - 3cost) after simplifying?

OpenStudy (theoreo):

No, I didn't simplify. I just put the answer in like that, and was using my graphing calculator with the unsimplified equation.

OpenStudy (anonymous):

well, you should simply so that it is less work when you solve

OpenStudy (theoreo):

Ok, I'll get to work on simplifying it. Might take me a bit though.

OpenStudy (anonymous):

I would suggest you do the second derivative again. Yours didn't match mime https://www.wolframalpha.com/input/?i=%5Cfrac%7B+2%5Csin%282t%29*%5Ccos%28t%29-%5Csin%28t%29*4%5Ccos%282t%29+%7D%7B+2%5Csin%282t%29%5E2*-%5Csin%282t%29+%7D

OpenStudy (theoreo):

Oh jeez, I missed a 2 in there and didn't notice this whole time. It should be \[\frac{ 2\sin(2t)*\cos(t)-\sin(t)*4\cos(2t) }{ 2\sin(t)^2-2\sin(2t) }\] Sorry for wasting all of yours and greenglasses' time trying to solve the wrong equation.

OpenStudy (anonymous):

lol, ok you found the mistake. But notice that sin(2t) = 2sint cost dy/dx = sint / (4sint cost) = (1/4)sect the second derivative is much easier to do if you simplify first derivative first

OpenStudy (theoreo):

So \[\frac{ 1 }{ 4\cos(t) } = 0\] = \[\frac{1}{\cos(t)} = 0\] Are there real solutions to this?

OpenStudy (anonymous):

also, forget about my simplified version earlier, it is also wrong lol

OpenStudy (anonymous):

(1/4)sect is only the first derivative. Second derivative is (1/4)sect / (-2sin(2t)) (1/4)sect / (-4sint cost) (1/4)csct sec^2(t)

OpenStudy (anonymous):

(1/4)csct sec^2(t) = 0 has no solution

OpenStudy (theoreo):

I forgot what that means for the concavity of the original curve.

OpenStudy (anonymous):

if the second derivative is positive, then it's concave up

OpenStudy (theoreo):

Ok, I see what you're saying. So the final answer in interval notation should be (0, pi)?

OpenStudy (theoreo):

But if the boundaries in the initial problem are 0 < t < pi ?